# Perfect Secret Key Generation for a class of Hypergraphical Sources

### v1

@@ II-B Hypergraphical Source Model
Let $s\in\mathbb{N}$ and let $Y=(Y_{1},Y_{2},\ldots,Y_{s})^{T}\sim\text{unif}\{\{0,1\}^{s}\}$. Let $M\in\{0,1\}^{l\times s}$ be a matrix whose rows we denote as $M_{1},M_{2},\ldots,M_{l}$. For any $1\leq l^{\prime}<l$ if $\text{span}\{M_{1},\ldots,M_{l^{\prime}}\}\cap\text{span}\{M_{l^{\prime}+1},\ldots,M_{l}\}=\{0\}$,5550 here denotes the vector $(\underbrace{0,0,\ldots,0}_{s\text{ times}})$. then $I(M_{1}Y,\ldots,M_{l^{\prime}}Y;M_{l^{\prime}+1}Y,\ldots,M_{l}Y)=0$. Furthermore, $H(M_{l^{\prime}+1}Y,\ldots,M_{l}Y)=\dim(\text{span}\{M_{l^{\prime}+1},\ldots,M_{l}\})$.
@@
Lemma 5 is now leveraged to derive lower bounds on the perfect secret key rate achievable through perfect omniscience obtained via linear communication.
Lemma 2.
Let $X_{\mathcal{M}}$ be a hypergraphical source defined on the underlying hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$. For a blocklength $n\in\mathbb{N}$ let there exists a linear communication scheme of $r<n|\mathcal{E}|$ bits achieving perfect omniscience. Then there exists a perfect secret key of rate at least $|\mathcal{E}|-\frac{r}{n}$.
@@
To see that claim holds, firstly observe that $K$ is perfectly recoverable as every party $i\in\mathcal{M}$ has access to $\xi_{\mathcal{E}}^{n}$ thanks to perfect omniscience. Lemma 5 shows that $I(K;F)=0$. Furthermore, as $\{M^{\prime}_{r^{\prime}+1},\ldots,M^{\prime}_{n|\mathcal{E}|}\}$ is linearly independent, Lemma 5 shows that $H(K)=n|\mathcal{E}|-r^{\prime}=\log|\mathcal{K}|$. Thus, the key $K$ is perfectly uniform, and hence $K$ is a perfect secret key. The rate of $K$ is thus $\frac{1}{n}H(K)=|\mathcal{E}|-\frac{r^{\prime}}{n}\geq|\mathcal{E}|-\frac{r}{n}$. ∎
We conclude this section by showing how the notion of packing can be used in tandem with the perfect secret key generation scheme of Lemma 2 to increase the perfect secret key rate available from hypergraphical sources.
A hypergraphical source $X_{\mathcal{M}}$ on the underlying hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$ is said to admit a packing for blocklength $n$, if there exists disjoint subsets $\mathcal{E}_{(1)},\mathcal{E}_{(2)},\ldots,\mathcal{E}_{(s)}\subseteq\mathcal{E}^{(n)}$ satisfying the constraint that for every $1\leq l\leq s$, the set $\mathcal{E}_{(l)}$ contains at most one single fictitious copy of a hyperedge $e\in\mathcal{E}$. The last constraint ensures that for every $1\leq l\leq s$, $\mathcal{H}_{(l)}:=(\mathcal{M},\mathcal{E}_{(l)})$ can be viewed as a hypergraph in its own right. Moreover, we define $X_{i_{(l)}}:=(\xi_{e(j)}:i\in e,e(j)\in\mathcal{E}_{(l)})$, and consequently $X_{\mathcal{M}_{(l)}}:=(X_{i_{(l)}}:i\in\mathcal{M})$ is a hypergraphical source.
Lemma 3.
Let $X_{\mathcal{M}}$ be a hypergraphical source on the underlying hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$. For a blocklength $n$, let $\mathcal{H}$ admits a packing, such that the hypergraphical source $X_{\mathcal{M}_{(l)}},1\leq l\leq s$, can generate a perfect secret key of $k_{l}$ bits using a blocklength of one. Then, $X_{\mathcal{M}}$ can generate a perfect secret key of rate $\frac{\sum_{l=1}^{s}k_{l}}{n}$.
Proof:
Let $K_{l},1\leq l\leq s$, be the perfect secret key of $k_{l}$ bits generated by $X_{\mathcal{M}_{(l)}}$ through public communication $F_{(l)}$. Thus, the overall public communication is $F=(F_{(l)}:1\leq l\leq s)$. Define $K=(K_{l}:1\leq l\leq s)$.
Now, observe that $\xi_{\mathcal{E}_{(1)}},\xi_{\mathcal{E}_{(2)}},\ldots,\xi_{\mathcal{E}_{(s)}}$ are mutually independent as $\mathcal{E}_{(1)},\mathcal{E}_{(2)},\ldots,\mathcal{E}_{(l)}$ are disjoint sets. As a consequence, $K$ is uniformly distributed, $I(K;F)=0$, and $H(K)=\sum_{l=1}^{s}k_{l}$. Hence, $K$ is the required perfect secret key, and its rate is $\frac{\sum_{l=1}^{s}k_{l}}{n}$. ∎
III Secret Key Construction Scheme for complete $t$-Uniform Hypergraphical Sources
In this section, we specialize the general hypergraphical model to the case where $\mathcal{H}$ is a complete $t$-uniform hypergraph, i.e., $\mathcal{E}=\binom{\mathcal{M}}{t}$. We denote the hyperedge variables by $\xi_{i_{1}i_{2}\ldots i_{t}}$ for $\{i_{1}i_{2}\ldots i_{t}\}\in\mathcal{E}$. We assume access to $t$ i.i.d. copies of the source, denoted $X_{\mathcal{M}}^{(1)},\dots,X_{\mathcal{M}}^{(t)}$. These $t$ copies constitute the necessary resources for the structural decomposition.
Definition 2 (Star Model $S_{m,i}$).
Consider a hypergraphical source defined on a complete $t$-uniform hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$. Fix a party $i\in\mathcal{M}$. The Star model $S_{m,i}$ is a pairwise source model defined on the vertex set $\mathcal{M}$ with edge set $\mathcal{E}_{i}=\big\{e^{\prime}\in\mathcal{E}\mid i\in e^{\prime}\big\}$. The random variable associated with an edge $e\in\mathcal{E}_{i}$ is the hyperedge variable $\xi_{e}$ from the original source.
Note that party $i$ observes all edge variables in the projected graph $S_{m,i}$ (since $i$ is a member of every hyperedge). Our construction proceeds in two phases:
  1. 1.
    Global Decomposition: We first define a bijection that partitions the aggregate resources from $t$ copies of the hypergraph into $m$ edge-disjoint Star models.
  2. 2.
    Local Protocol Execution: Once the Star models are defined, each party $i\in\mathcal{M}$ independently executes the secret key generation protocol on the specific Star model $S_{m,i}$ assigned to it.
{restatable}
[Global Source Decomposition]lemmasourcedecomp Given $t$ independent copies of a complete $t$-uniform hypergraphical source $\mathcal{H}=(\mathcal{M},\mathcal{E})$, the aggregate collection of hyperedge copies can be perfectly partitioned into $m$ edge-disjoint Star models $\{S_{m,1},S_{m,2},\dots,S_{m,m}\}$, where $S_{m,i}$ is anchored at party $i$.
1
2
Input: Observations from $t$ independent copies of the source: $X_{\mathcal{M}}^{(1)},\dots,X_{\mathcal{M}}^{(t)}$.
Output: Public communication tuple $F=(F^{(1)},F^{(2)},\dots,F^{(m)})$.
3
Init: Initialize empty communication lists $F^{(i)}\leftarrow\emptyset$ for all $i\in\mathcal{M}$.
4
5foreach hyperedge $e\in\mathcal{E}$ do
6   Fix an ordering of the vertices of $e$ as $(u_{1},u_{2},\ldots,u_{k})$, $k\in\{1,\ldots,t\}$.
7   Assign the $k$th copy of the hyperedge $e^{(k)}$ to the Star model $S_{m,u_{k}}$ anchored at $u_{k}$.
8
9
10foreach party $i\in\mathcal{M}$ do
11   Let $\mathcal{E}_{i}$ be the set of edge variables assigned to $S_{m,i}$.
12   Choose an arbitrary helper party $j\in\mathcal{M}\setminus\{i\}$. Let $\mathcal{E}_{ij}=\{e\in\mathcal{E}\mid i\in e,j\in e\}$.
13
14 foreach $e\in\mathcal{E}_{i}\setminus\mathcal{E}_{ij}$ do
15    $M_{\text{xor}}\leftarrow 0$.
16
[…]
18
19       $M_{\text{xor}}\leftarrow M_{\text{xor}}\oplus\xi_{\{i,j\}\cup A}$.
20
21    $msg\leftarrow\xi_{e}\oplus M_{\text{xor}}$.
22      Append msgmsg to $F^{(i)}$.
23
24  Party $i$ sends $F^{(i)}$.
Algorithm 1 Global Secret Key Construction Protocol
In Algorithm 1, the global execution consists of $m$ independent instances of a Star protocol. For a specific party $i$ acting as the anchor for $S_{m,i}$, the number of messages sent is $\binom{m-1}{t-1}-\binom{m-2}{t-2}$. We define the secret key $K^{(i)}$ generated by this specific instance as the subset of edges in $S_{m,i}$ known to both $i$ and its helper $j$. By the construction of the Star model, this set $K^{(i)}$ corresponds to the variables in $\mathcal{E}_{ij}$, which has cardinality $\binom{m-2}{t-2}$. The subsequent lemmas formally prove that for any such instance, $K^{(i)}$ is perfectly recoverable by all parties and perfectly secret from the public communication.
{restatable}
[Perfect Recoverability]lemmaperfectrec For a fixed anchor $i$, after the execution of Algorithm 1, every party $v\in\mathcal{M}$ can perfectly recover the secret key $K^{(i)}$.
{restatable}
[Perfect Secrecy and Uniformity]lemmaperfectsec The secret key $K^{(i)}$ generated by Algorithm 1 is uniformly distributed and perfectly secret with respect to the public communication $F$. That is, $H(K^{(i)})=\log|\mathcal{K}|$ and $I(K^{(i)};F)=0$.
Theorem 4 (Main Result).
The protocol defined by Algorithm 1 constitutes a capacity-achieving perfect secret key generation scheme for complete $t$-uniform hypergraphical sources.
Proof:
Our protocol begins with the Global Decomposition phase described in Lemma 2. We utilize $t$ copies of the source to construct exactly $m$ edge-disjoint Star models $S_{m,1},\dots,S_{m,m}$, where $S_{m,i}$ is anchored at party $i$. This decomposition is deterministic and exhaustive, consuming all edge resources without collision. We then assume, without loss of generality, that the indices $\{1,\dots,m\}$ assigned to the Star models correspond to the party indices. We execute Algorithm 1 on each of the $m$ Star models in parallel.
For the instance $S_{m,i}$, party $i$ acts as the sender and selects an arbitrary helper $j_{i}\in\mathcal{M}\setminus\{i\}$. From Lemmas 1 and 1, each instance generates a key $K^{(i)}$ that is perfectly recoverable, perfectly secret, and uniform. Since the Star models are edge-disjoint (operating on distinct copies of variables), the execution of the $m$ protocols is mutually independent. Thus, the aggregate properties hold for the combined key.
To calculate the rate, we sum the key bits generated by each model. For a Star model anchored at $i$, the key size is equal to the number of shared edges with helper $j_{i}$, denoted $|\mathcal{E}_{ij}|$. In a complete $t$-uniform hypergraph, for any pair of parties $i,j$, the number of edges containing both is exactly $\binom{m-2}{t-2}$. Thus, $|K^{(i)}|=\binom{m-2}{t-2}$ for all $i$. Since there are $m$ such models, the total number of key bits generated is $m\times\binom{m-2}{t-2}$. The rate is normalized by the $t$ copies of the source utilized in the decomposition $R_{\mathrm{SK}}=\frac{m}{t}\binom{m-2}{t-2}$. Finally, we verify optimality. The secret key capacity for complete $t$-uniform hypergraphs is given as $C_{\mathrm{SK}}(X_{\mathcal{M}})=\frac{t-1}{m-1}\binom{m}{t}=\frac{m}{t}\binom{m-2}{t-2}$. Since $R_{\mathrm{SK}}=C_{\mathrm{SK}}(X_{\mathcal{M}})$, the scheme is capacity achieving. ∎
IV Secret Key Construction Scheme for $3$-uniform Hypergraphs
In this section, we provide a new secret key generation scheme for $3$-uniform hypergraphs (not necessarily complete), and we identify a subclass where our scheme achieves capacity. We start by defining the notion of an induced cycle 33-uniform hypergraph, which is used as a key construct in our scheme.
Definition 3 (Induced cycle 33-uniform hypergraphical source).
A $3$-uniform hypergraphical source of the form $\mathcal{H}_{i}=\{\mathcal{M},\mathcal{E}\}$ is said to be an induced cycle 33-uniform hypergraphical source if: $i\in e\ \forall\ e\in\mathcal{E}$ and the projected graph from ii [^6] For a hypergraph $\mathcal{H}=\{\mathcal{M},\mathcal{E}\}$, the projected graph from $i\in\mathcal{M}$ is the graph $G=\{\mathcal{M}\setminus\{i\},\mathcal{E}_{i}\}$ where $\mathcal{E}_{i}=\{e\setminus\{i\}:e\in\mathcal{E},i\in e\}$. (called the anchor vertex) forms a cycle.
{restatable}
[Secret Key Capacity for the induced cycle 33-uniform hypergraph]propositionprop There exists a scheme for the induced cycle 33-uniform hypergraph that generates $2$ secret key bits, depending on the parity of $m$. Let k=m1k=m-1, then the scheme is given as follows:
  1. 1.
    If $k=3l$ or $k=3l+2$ the anchor $i$ sends the following k2k-2 messages: $e_{1}\oplus e_{2}\oplus e_{3}$, $e_{2}\oplus e_{3}\oplus e_{4}$, … $e_{k-2}\oplus e_{k-1}\oplus e_{k}$.
  2. 2.
    If $k=3l+1$, the anchor $i$ sends the messages $e_{1}\oplus e_{2}\oplus e_{3}$, …, $e_{k-3}\oplus e_{k-2}\oplus e_{k-1}$, and finally the message $e_{k-1}\oplus e_{k}\oplus e_{1}$.
Now, we describe the generic scheme for $3$-uniform hypergraphs.
Definition 4 (Generic Scheme for $3$-uniform hypergraphs).
Consider a 33-uniform hypergraph $\mathcal{H}=\{\mathcal{M},\mathcal{E}\}$.
  1. 1.
    Fix a set of anchor points $A\subseteq\mathcal{M}$.
  2. 2.
    For each anchor $i\in A$, suppose $k_{i}$ copies of the projected graph from $i$ admit a Hamiltonian decomposition of size $p_{i}$.
  3. 3.
    Each edge $e$ is used $t_{e}$ times where $t_{e}=\sum_{i\in A}k_{i}\mathbf{1}\{i\in e\}$, choose $n=\max_{e\in\mathcal{E}}t_{e}$.
  4. 4.
    Consider $n$ independent copies of the original source $\mathcal{H}$ and for each anchor point $i$, consider the $p_{i}$ Hamiltonian decompositions and observe that each decomposition corresponds to an induced cycle hypergraph. We can now run the scheme described in Proposition 3 as a subroutine on each of these.
Observe that since each run of the subroutine gives us $2$ bits of secret key (see Proposition 3), we get a scheme with rate $\frac{2}{n}\sum_{i\in A}p_{i}$. We now identify three different classes of $3$-uniform hypergraphs for which the scheme described in Defintion 4 is capacity achieving.
First we consider the class of complete $3$-uniform hypergraph on $m$ vertices. {restatable}corollarytypea For the complete $3$-uniform hypergraph $K_{m,3}$ on $m$ vertices, applying the generic scheme (see Definition 4) with the anchor set $A=[m]$ achieves capacity.
Next, we consider a particular family of $3$-uniform hypergraphs whose projected graphs admit Hamiltonian decompositions. An example of such a hypergraph are hypergraphs wherein the projected graph from a vertex is a Paley graph [ALSPACH2012113].
{restatable}
corollarytypeb Consider a hypergraphical source $\mathcal{H}=\{\mathcal{M},\mathcal{E}\}$ constructed such that there exists a specific vertex $i\in\mathcal{M}$ where $i\in e$ for all $e\in\mathcal{E}$, such that $k$ independent copies of the projected graph from $i$ admit a Hamiltonian cycle packing of size $p$. Applying the generic scheme with the $i$ as the sole anchor achieves capacity. Further, we show that such a hypergraphical source is Type-S.
Finally, we identify a class of $3$-uniform hypergraphs that we term Hollow 3D Kite hypergraphs. We show that the generic scheme achieves capacity and such a hypergraphical source is Type-S.
{restatable}
corollarytypec Consider a hypergraphical source with $m=4r+1$ vertices. Let the vertex set be partitioned into two sets: $V_{1}=\{1,\dots,2r+1\}$ and $V_{2}=\{2r+2,\dots,4r+1\}$. The edge set is defined as $\mathcal{E}=\left\{\{i,j,k\}:\{i,j\}\subseteq V_{1},k\in V_{2}\right\}$. The generic scheme applied with the anchor set $A=V_{1}$ achieves capacity. Further, such a hypergraphical source is Type-S.
Appendix A Ommitted Proofs
A-A Omitted Proofs of Section III
@@
Note that exactly $t$ vertices participate in every hyperedge $e\in\mathcal{E}$, and there are $t$ independent copies of every hyperedge. Fix an arbitrary ordering of vertices in every hyperedge. For each hyperedge, assign the $k$th copy of hyperedge $e$ to the $k$th vertex. Now consider any vertex $v\in\mathcal{M}$. The vertex $v$ receives exactly one copy of every hyperedge it is a part of, and hence is assigned the edge set $\mathcal{E}_{v}$. This is exactly the star model $S_{m,v}$. Since the choice of vertex was arbitrary, we have $m$ different star models each anchored at a different vertex. ∎
@@
Let $j$ be the helper vertex fixed in Algorithm 1, so the secret key is $K^{(i)}=\mathcal{E}_{ij}$. We analyze the recoverability for three distinct cases covering all parties in $\mathcal{M}$.
  1. 1.
    Case 1: The Anchor Party $i$. Party $i$ is a member of every hyperedge in the Star model $S_{m,i}$. Consequently, $i$ directly observes every variable $\xi_{e}$ for all $e\in\mathcal{E}_{i}$, including the subset $\mathcal{E}_{ij}$. Thus, $i$ possesses $K^{(i)}$ directly from its source observations.
  2. 2.
    Case 2: The Helper Party $j$. The secret key is defined as the set of edges $\mathcal{E}_{ij}=\{e\in\mathcal{E}\mid i\in e,j\in e\}$. By definition, party $j$ is a member of every hyperedge in this set. Therefore, $j$ directly observes every variable $\xi_{e}\in K^{(i)}$. (Note: Although $j$ also observes the public messages $M_{e}$, it does not need them to recover the key; it uses the key and the messages to recover the remaining variables in $\mathcal{E}_{i}\setminus\mathcal{E}_{ij}$, effectively reconstructing the entire Star).
  3. 3.
    Case 3: Any Other Party $k\in\mathcal{M}\setminus\{i,j\}$. We show that $k$ can recover any specific key bit $\xi_{E^{*}}\in K^{(i)}$, where $E^{*}\in\mathcal{E}_{ij}$. If $k\in E^{*}$, then $k$ observes $\xi_{E^{*}}$ directly. If $\notin E^{*}, kk$ must recover $\xi_{E^{*}}$ from the public communication. The algorithm transmits a message $M_{e}$ for the hyperedge $e=(E^{*}\setminus\{j\})\cup\{k\}$. Note that $e\in\mathcal{E}_{i}\setminus\mathcal{E}_{ij}$ because it contains $i$ (from $E^{*}$) and $k$, but explicitly excludes $j$. Since $k\in e$, party $k$ knows variable $\xi_{e}$. The message structure is:
    $M_{e}=\xi_{e}\oplus\bigoplus_{A\subseteq e\setminus\{i\},|A|=t-2}\xi_{i,j,A}.$
    Party $k$ computes the XOR sum. The sum runs over subsets $A$ of $e\setminus\{i\}$. The set $e\setminus\{i\}$ is exactly $(E^{*}\setminus\{i,j\})\cup\{k\}$. If the subset $A$ is formed by excluding $k$, then $A=E^{*}\setminus\{i,j\}$. The corresponding term is $\xi_{i,j,A}=\xi_{E^{*}}$. This is the target variable $k$ wishes to recover. On the other hand, if the subset $A$ includes $k$, then the corresponding edge $\{i,j\}\cup A$ contains $k$. As a member of this hyperedge, $k$ observes the variable directly.
    Thus, the message $M_{e}$ relates the unknown $\xi_{E^{*}}$ to values known by $k$ (the message $M_{e}$, $\xi_{e}$, and all other terms in the sum). Party $k$ can therefore solve for $\xi_{E^{*}}$. Since this holds for all $E^{*}\in\mathcal{E}_{ij}$, $k$ recovers the full secret key $K^{(i)}$. Consequently, having recovered the key, $k$ can also decode all messages $M_{e}$ to recover the entire Star $\mathcal{E}_{i}$, matching the knowledge of $j$.

### v2

@@ Appendix A Omitted proofs of Section II-B
Let $s\in\mathbb{N}$ and let $Y=(Y_{1},Y_{2},\ldots,Y_{s})^{T}\sim\text{unif}\{\{0,1\}^{s}\}$. Let $M\in\{0,1\}^{l\times s}$ be a matrix whose rows we denote as $M_{1},M_{2},\ldots,M_{l}$. For any $1\leq l^{\prime}<l$ if $\text{span}\{M_{1},\ldots,M_{l^{\prime}}\}\cap\text{span}\{M_{l^{\prime}+1},\ldots,M_{l}\}=\{0\}$,1010100 here denotes the vector $(\underbrace{0,0,\ldots,0}_{s\text{ times}})$. then $I(M_{1}Y,\ldots,M_{l^{\prime}}Y;M_{l^{\prime}+1}Y,\ldots,M_{l}Y)=0$. Furthermore, $H(M_{l^{\prime}+1}Y,\ldots,M_{l}Y)=\dim(\text{span}\{M_{l^{\prime}+1},\ldots,M_{l}\})$.
@@
Lemma 10 is now leveraged to prove Lemma II-B.
[…]
@@
To see that the claim holds, firstly observe that $K$ is perfectly recoverable as every party $i\in\mathcal{M}$ has access to $\xi_{\mathcal{E}}^{n}$ thanks to perfect omniscience. Lemma 10 shows that $I(K;F)=0$. Furthermore, as $\{M^{\prime}_{r^{\prime}+1},\ldots,M^{\prime}_{n|\mathcal{E}|}\}$ is linearly independent, Lemma 10 shows that $H(K)=n|\mathcal{E}|-r^{\prime}=\log|\mathcal{K}|$. Thus, the key $K$ is perfectly uniform, and hence $K$ is a perfect secret key. The rate of $K$ is thus $\frac{1}{n}H(K)=|\mathcal{E}|-\frac{r^{\prime}}{n}\geq|\mathcal{E}|-\frac{r}{n}$. ∎
@@
Let $K_{l},1\leq l\leq s$, be the perfect secret key of $k_{l}$ bits generated by $X_{\mathcal{M}_{(l)}}$ through public communication $F_{(l)}$. Then, $H(K_{l})=k_{l}$ and $I(K_{l};F_{(l)})=0$. Denote the overall public communication by $F=(F_{(l)}:1\leq l\leq s)$ and define $K=(K_{l}:1\leq l\leq s)$. We claim that $K$ is the required perfect secret key.
To prove the claim, firstly observe that $K$ is perfectly recoverable, since the individual $K_{l}$s are perfectly recoverable. Next, note that $\xi_{\mathcal{E}_{(1)}},\xi_{\mathcal{E}_{(2)}},\ldots,\xi_{\mathcal{E}_{(s)}}$ are mutually independent as $\mathcal{E}_{(1)},\mathcal{E}_{(2)},\ldots,\mathcal{E}_{(l)}$ are disjoint sets. As a consequence, $K_{1},\ldots,K_{s}$ are mutually independent, and so are $F_{(1)},F_{(2)},\ldots,F_{(l)}$. Hence, $H(K)=\sum_{l=1}^{s}H(K_{l})=\sum_{l=1}^{s}k_{l}$, and so $K$ is uniformly distributed over the key alphabet $\{0,1\}^{\sum_{l=1}^{s}k_{l}}$. Finally, the mutual independece of $K_{l}$s and $F_{(l)}$s coupled with the fact that $I(K_{l};F_{(l)})=0$ for all $1\leq l\leq s$ ensures that $I(K;F)=0$, and hence $K$ is perfectly secure. The rate of $K$ is thus $\frac{1}{n}H(K)=\frac{1}{n}\sum_{l=1}^{s}k_{l}$ as required. ∎
Appendix B Omitted proofs of Section III
@@
We obtain the packing as follows. Fix any $e\in\mathcal{E}$ and order its vertices in the usual ascending order. In other words, let $e=\{v_{1},v_{2},\ldots,v_{t}\}$ where $v_{1}<v_{2}<\ldots v_{t}$. Then assign $e(l)$ to $\mathcal{E}_{(v_{l})}$. For example, if $m=4,t=3,$ and $e=\{1,2,4\}$, then $e(1)$ is assigned to $\mathcal{E}_{(1)}$, $e(2)$ is assigned to $\mathcal{E}_{(2)}$, and $e(3)$ is assigned to $\mathcal{E}_{(4)}$. Observe that by construction $\mathcal{E}_{(1)},\mathcal{E}_{(2)},\ldots,\mathcal{E}_{(m)}$ is indeed a packing as any $e(j),e\in\mathcal{E},1\leq j\leq t$ is assigned to exactly one $\mathcal{E}_{(l)}$. It remains to show that for any $1\leq l\leq m$, the set $\mathcal{E}_{(l)}$ consists of exactly one fictitious copy of every edge in $\mathcal{E}_{l}$.
To proceed, fix any $l\in\mathcal{M}$ and $e\in\mathcal{E}_{l}$. It is enough to show that for exactly one $1\leq j\leq t$, $e(j)\in\mathcal{E}_{(l)}$. Now, let $e=\{v_{1},v_{2},\ldots,v_{t}\}$. Then, as $l\in e$, there exists exactly one $1\leq j\leq t$ such that $v_{j}=l$. Thus, by construction, only $e(j)$ gets assigned to $\mathcal{E}_{(l)}$ as required. ∎
Appendix C Omitted Proofs of Section IV